# Context-Aware Regularization with Markovian Integration for Attention-Based Nucleotide Analysis

### v1

@@ Abstract
Transformers have revolutionized nucleotide sequence analysis, yet capturing long-range dependencies remains challenging. Recent studies show that autoregressive transformers often exhibit Markovian behavior by relying on fixed-length context windows for next-token prediction. However, standard self-attention mechanisms are computationally inefficient for long sequences due to their quadratic complexity and do not explicitly enforce global transition consistency.
@@
We evaluate CARMANIA across diverse genomic tasks, including regulatory element prediction, functional gene classification, taxonomic inference, antimicrobial resistance detection, and biosynthetic gene cluster classification. CARMANIA outperforms the previous best long-context model by at least 7%, matches state-of-the-art on shorter sequences (exceeding prior results on 20/40 tasks while running $\sim$2.5$\times$ faster), and shows particularly strong improvements on enhancer and housekeeping gene classification tasks—including up to a 34% absolute gain in Matthews correlation coefficient (MCC) for enhancer prediction. The TM loss boosts accuracy in 33 of 40 tasks, especially where local motifs or regulatory patterns drive prediction. This enables more effective modeling of sequence-dependent biological features while maintaining robustness across non-coding and low-signal regions.Code available at https://github.com/EESI/carmania.
@@ 1 Introduction
Despite advances in sequencing, analyzing large-scale biological data remains challenging. Traditional methods—such as motif detection, sequence alignment, and Markov models—capture short-range dependencies well Delcher et al. (2007); Steinegger and Söding (2017); Wood et al. (2019), but often fail to represent complex or long-range patterns, especially in diverse or non-canonical sequences. Recent work has applied NLP techniques to biological sequences Nguyen et al. (2024); Cahyawijaya et al. (2022); Schiff et al. (2024), with large language models (LLMs) showing strong ability to capture complex dependencies. However, genomic sequences are often extremely long, making full attention computationally infeasible. To address this, transformer models use mechanisms like sliding window and sparse attention Dong et al. (2025); Zaheer et al. (2020), which expand the receptive field but may still struggle to capture long-range dependencies effectively in genomics.
@@
Our approach, named CARMANIA (Context-Aware Regularization with Markovian Integration for Attention-Based Nucleotide Analysis), supplements the standard next-token prediction objective with an auxiliary loss. This loss aligns the model’s predicted first-order transition matrix with one derived from the full input sequence, encouraging consistency with its empirical token transition patterns. By doing so, CARMANIA captures the probabilistic structure of genomic sequences and improves context-awareness during next-token prediction. This concept is illustrated in Figure 1. In addition to improving context modeling, our pretraining approach supports efficient domain adaptation, as conserved transition patterns across species enable generalization with minimal fine-tuning Bergeron et al. (2023).
@@ 3 Pretraining Long Sequences With Markovian Knowledge
We adopt a causal Transformer inspired by LLaMATouvron et al. (2023), where each layer computes hidden representations $\mathbf{h}^{(l)}$ from input tokens $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$ and the previous layer’s outputs $\mathbf{h}^{(l-1)}$, using multi-head self-attention to predict the next token. To scale to long sequences, we replace full attention with Sliding Window Attention (SWA), where each token attends only to a fixed-size window of preceding tokens, ${x_{t-w},\dots,x_{t-1}}$, reducing complexity from $O(n^{2})$ to $O(n)$. We further improve efficiency with FlashAttention-2 Dao (2023) and enhance positional encoding via Rotary Positional Embeddings (RoPE) Su et al. (2024), enabling better extrapolation to long contexts.
@@ 3.1 Unified $n$-th Order Transition Tensor Formulation
To complement the next-token prediction objective, we introduce a unified probabilistic framework that captures higher-order dependencies through a transition tensor. Let $\mathbf{P}_{t}\in\mathbb{R}^{V}$ denote the model’s predicted probability distribution over the vocabulary $\mathcal{V}$ at position $t$, where $[\mathbf{P}_{t}]_{i}=\mathbb{P}(x_{t}=i\mid x_{<t})$.
@@ 4 Experiments
Table 1: Ablation study showing the impact of TM loss, attention type, and sequence length on F1, BLEU, and compute cost.
@@ 4.2 Ablation Studies
Higher-Order Transition Matrices. As shown in Supplementary Tables 8 and 9, applying second-order TM loss led to performance degradation across tasks compared to first-order TM or no TM. This drop is likely due to the sparsity of higher-order transitions in biological sequences, which makes the loss signal unstable. In contrast, first-order TM loss consistently improves performance by capturing reliable and biologically meaningful pairwise dependencies.
@@ 4.4 Downstream Task Evaluations
As shown in Table 3, CARMANIA achieves top-1 performance on 5 out of 18 tasks and outperforms the average of all baselines by over 3% across metrics. Notably, in enhancer classification tasks, CARMANIA achieves up to 34% absolute improvement in MCC over previous models Zhou et al. (2023); Avsec et al. (2021); Schiff et al. (2024); Nguyen et al. (2024); Dalla-Torre et al. (2024). These results demonstrate the model’s strong ability to capture regulatory signals from sequence. A notable gain is observed on the H4ac histone mark, where adding the TM loss yields over 40% relative improvement in MCC. This likely reflects the clustered nature and sequence specificity of acetylation sites in the H4 N-terminal tail Struhl (1998); Ma et al. (2024). In contrast, marks like H3K36me3 and H3K79me3, which are more spatially diffuse and functionally broad Jha and Strahl (2014); Struhl (1998), show limited improvement, likely due to weaker dependence on local sequence patterns(as highlighted in the Table).
@@ Appendix A Supplementary Material
A.1 Domain Adaptation and Generalization
We evaluated several pre-trained genomic models on the AMR dataset without fine-tuning. The dataset, derived from MEGARes Bonin et al. [2023] and CARD Jia et al. [2016], includes three classification tasks: Gene Family, Resistance Mechanism, and Drug Class Yoo et al. [2024]. Labels were assigned based on the closest match (best hit) in embedding space using FAISSDouze et al. [2024] on frozen representations.
As shown in Table 6, CARMANIA achieves the highest F1 Macro scores across all tasks, indicating that it produces more informative embeddings for AMR classification than prior models, including HyenaDNANguyen et al. [2024], CaduceusSchiff et al. [2024], MetaBERTa Refahi et al. [2025], and Nucleotide TransformersDalla-Torre et al. [2024].
Model GeneFamily Resist-Mech DrugClass
MetaBERTa (35.2M) 0.650 0.898 0.886
HyenaDNA (1.6M) 0.623 0.840 0.880
Caduceus-PH (1.9M) 0.597 0.823 0.839
NucleotideTrans (2.5B) 0.611 0.856 0.859
w/o TM (83M) 0.728 0.974 0.931
CARMANIA (83M) 0.733 0.975 0.942
A.2 Effect of TM Loss on Convolution-Based Model HyenaDNA
@@
A.3 Sensitivity Analysis: Effect of $\beta$
@@
A.4 Higher-Order Transition Matrix
We explored the impact of incorporating TM loss at different Markov orders in CARMANIA, with results shown in Tables 8 and 9. Surprisingly, while the first-order TM consistently improves downstream classification performance across tasks, the second-order TM performs worse than both the first-order TM and the baseline without TM.
@@ A.4 Higher-Order Transition Matrix
A.5 Wide vs. Deep: Architectural Trade-offs for DNA Modeling
@@
A.6 Model Size Scalability
@@
A.7 Impact of Sequence Length Scaling
@@
A.8 Pre-training Dataset
@@ A.9 Downstream Datasets
Scorpio-Gene-Taxa Refahi et al. [2025] 6 Multi-class (497–1,929) 114–13,227
Antimicrobial Resistance Prediction Yoo et al. [2024] 3 Multi-class (497–1,929) 211–5,274

### v2

@@ Abstract
Transformers have revolutionized nucleotide sequence analysis, yet capturing long‑range dependencies remains challenging. Recent studies show that autoregressive transformers often exhibit Markovian behavior by relying on fixed-length context windows for next-token prediction. However, standard self-attention mechanisms are computationally inefficient for long sequences due to their quadratic complexity and do not explicitly enforce global transition consistency.
@@
[…]
@@ 1 Introduction
Despite advances in sequencing, analyzing large-scale biological data remains challenging. Traditional methods—such as motif detection, sequence alignment, and Markov models—capture short-range dependencies well Delcher et al. (2007); Steinegger and Söding (2017); Wood et al. (2019), but often fail to represent complex or long-range patterns, especially in diverse or non-canonical sequences. Recent work has applied NLP techniques to biological sequences Nguyen et al. (2024); Cahyawijaya et al. (2022); Schiff et al. (2024); Bo et al. (2025), with large language models (LLMs) showing strong ability to capture complex dependencies. However, genomic sequences are often extremely long, making full attention computationally infeasible. To address this, transformer models use mechanisms like sliding window and sparse attention Dong et al. (2025); Zaheer et al. (2020), which expand the receptive field but may still struggle to capture long-range dependencies effectively in genomics.
@@
Our approach, named CARMANIA (Context-Aware Regularization with Markovian Integration for Attention-Based Nucleotide Analysis), supplements the standard next-token prediction objective with an auxiliary loss. This loss aligns the model’s predicted first-order transition matrix with one derived from the full input sequence, encouraging consistency with its empirical token transition patterns. By doing so, CARMANIA captures the probabilistic structure of genomic sequences and improves context-awareness during next-token prediction.
To scale pretraining to long genomic sequences, we replace full attention with sliding-window attention (SWA), a sparse mechanism that limits each token’s receptive field to a fixed-size window of preceding tokens, reducing complexity from $O(n^{2})$ to $O(kn)$ ($k\ll n$) while preserving biologically meaningful local interactions. TM loss complements this by reinforcing short-range dependencies within the local window while promoting global consistency, since the transition matrix is computed over the entire input sequence. Together, these components enable a balance between local motif recognition and broader statistical alignment. In practice, CARMANIA handles effective contexts up to 160 kbp, which makes it, to our knowledge, the longest-context transformer-based genomic language model to date. An overview of our framework is illustrated in Figure 1.
In addition to improving context modeling, our pretraining approach supports efficient domain adaptation, as conserved transition patterns across species enable generalization with minimal fine-tuning Bergeron et al. (2023).
@@ 3 Pretraining Long Sequences With Markovian Knowledge
We adopt a causal Transformer inspired by LLaMATouvron et al. (2023), where each layer computes hidden representations $\mathbf{h}^{(l)}$ from input tokens $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$ and the previous layer’s outputs $\mathbf{h}^{(l-1)}$, using multi-head self-attention to predict the next token. To scale to long sequences, we replace full self-attention with SWA, where each token attends only to a fixed-size $k$ window of preceding tokens, $x_{t-k},\dots,x_{t-1}$}, reducing complexity from $O(n^{2})$ to $O(kn)$. This not only improves efficiency but also reflects the local nature of most biological dependencies in genomic sequences. We further improve efficiency with FlashAttention-2 Dao (2023) and enhance positional encoding via Rotary Positional Embeddings (RoPE) Su et al. (2024), enabling better extrapolation to long contexts.
@@ 3.1 Unified $n$-th Order Transition Tensor Formulation
To complement the next-token prediction objective, we introduce a unified probabilistic framework that captures higher-order dependencies through a transition tensor. Let $\mathbf{P}_{t}\in\mathbb{R}^{V}$ denote the model’s predicted probability distribution over the vocabulary $\mathcal{V}$ at position $t$, where $\mathbf{P}_{t}]_{i}=\mathbb{P}(x_{t}=i\mid x_{<t})$.
@@ 4 Experiments
Table 1: Ablation study showing the impact of TM loss and attention type on model performance (F1 Macro, BLEU) and compute cost. All models are trained on the Scorpio-Gene–Taxa dataset. F1 Macro is evaluated on the Antimicrobial Resistance (AMR) classification tasks, while BLEU and Relative FLOPs are computed on the Human Genome–Short dataset using 10,000 fragments from the GRCh38/hg38 reference genome. Relative FLOPs are reported with respect to the Sliding Window attention baseline, which has an absolute compute cost of approximately $16.7\times 10^{12}$ FLOPs per sequence.
@@ 4.2 Ablation Studies
Higher-Order Transition Matrices. As shown in Supplementary Tables 10 and 11, applying second-order TM loss led to performance degradation across tasks compared to first-order TM or no TM. This drop is likely due to the sparsity of higher-order transitions in biological sequences, which makes the loss signal unstable. In contrast, first-order TM loss consistently improves performance by capturing reliable and biologically meaningful pairwise dependencies.
@@ 4.4 Downstream Task Evaluations
As shown in Table 3, CARMANIA achieves top-1 performance on 5 out of 18 tasks and outperforms the average of all baselines by over 3% across metrics. Notably, in enhancer classification tasks, CARMANIA achieves up to 34% absolute improvement in MCC over previous models Zhou et al. (2023); Avsec et al. (2021); Schiff et al. (2024); Nguyen et al. (2024); Dalla-Torre et al. (2024). These results demonstrate the model’s strong ability to capture regulatory signals from sequence. A notable gain is observed on the H4ac histone mark, where adding the TM loss yields over 40.9% relative improvement in MCC. The benefit of the TM loss, however, is not uniform across all genomics tasks. In enhancer prediction, where positive and negative classes differ strongly in their local motif distributions (KL divergence = 9.28), the TM loss provides no measurable improvement ($\Delta=0.0\%$). In splice-site donor prediction (KL = 0.99), it yields a modest 5.2% gain, and in the H4ac histone mark dataset (KL = 0.64), a substantial 40.9% improvement. These results demonstrate that TM-driven performance gains inversely track inter-class bigram separability, confirming that TM loss is most effective when local sequence features are ambiguous and long-range dependencies become critical. Biologically, this aligns with our understanding that enhancer regions possess clear, unique motifs that NT loss can easily learn, whereas splice sites rely on short, dispersed sequence cues. In contrast, H4 acetylation involves multiple lysine sites (K5, K8, K12, K16) on the H4 N-terminal tail that act cooperatively to regulate chromatin accessibility Dion et al. (2005); Ma et al. (2024). Instead of strong individual motifs, these sites follow diffuse and combinatorial sequence patterns, making TM’s global co-occurrence modeling particularly advantageous.
@@
Domain Adaptation and Generalization
We evaluated several pre-trained genomic models on the AMR dataset without fine-tuning. The dataset, derived from MEGARes Bonin et al. (2023) and CARD Jia et al. (2016), includes three classification tasks: Gene Family, Resistance Mechanism, and Drug Class Yoo et al. (2024). Labels were assigned based on the closest match (best hit) in embedding space using FAISSDouze et al. (2024) on frozen representations.
As shown in Table 5, CARMANIA achieves the highest F1 Macro scores across all tasks, indicating that it produces more informative embeddings for AMR classification than prior models, including HyenaDNANguyen et al. (2024), CaduceusSchiff et al. (2024), MetaBERTa Refahi et al. (2025), and Nucleotide TransformersDalla-Torre et al. (2024).
Model GeneFamily Resist-Mech DrugClass
MetaBERTa (35.2M) 0.650 0.898 0.886
HyenaDNA (1.6M) 0.623 0.840 0.880
Caduceus-PH (1.9M) 0.597 0.823 0.839
NucleotideTrans (2.5B) 0.611 0.856 0.859
w/o TM (83M) 0.728 0.974 0.931
CARMANIA (83M) 0.733 0.975 0.942
@@ Appendix A Supplementary Material
A.1 Effect of TM Loss on Other Domains: Protein Modeling
While CARMANIA was primarily developed for genomic sequences, we further examined the generality of the proposed TM loss in other biological domains. In theory, TM loss should benefit any modeling task that relies on capturing long-range dependencies. To validate this hypothesis, we adapted the CARMANIA architecture for amino acid sequences and trained it on the Scorpio-Gene–Taxa protein dataset. Specifically, we constructed a $20\times 20$ bigram frequency matrix representing pairwise transition statistics between standard amino acids and incorporated TM loss into the training objective.
As summarized in Table 7, incorporating TM loss consistently accelerated convergence and improved predictive accuracy across all hierarchical levels. These results confirm that TM loss generalizes beyond nucleotide modeling and effectively enhances protein-sequence representations.
Table 7: Performance comparison on the Scorpio-Gene–Taxa dataset using protein sequences.
Test Set Gene-Heldout Taxa-Heldout
Model Phylum Class Order Family Gene Phylum Class Order Family Gene
Without TM Loss 91.1 84.9 70.6 51.3 99.5 19.8 10.9 3.3 1.1 98.4
With TM Loss 91.4 85.5 71.0 51.5 99.6 20.5 11.6 3.8 1.5 98.7
A.2 Effect of TM Loss on Convolution-Based Architecture: HyenaDNA
@@
A.3 Effect of TM Loss on Mamba-Based Architecture: Caduceus
Caduceus Schiff et al. [2024] is a bi-directional, reverse-complement equivariant DNA language model built upon the MambaDNA architecture, which leverages selective state-space modeling for long-range sequence processing. While Caduceus is pretrained using a masked language modeling (MLM) objective distinct from our causal next-token prediction (NT) setup we incorporated the proposed TM loss into the MLM objective while keeping all other configurations unchanged. This modification allows us to assess whether TM loss provides additional benefits beyond the CARMANIA backbone.
For this experiment, we used the largest available pretrained variant of Caduceus (approximately 7.7M parameters) and fine-tuned it on the Scorpio Gene–Taxa dataset with and without TM loss. Incorporating TM loss led to a notable performance boost, particularly on the gene classification task, yielding a +12% gain on the test set and +6% on the held-out set. These results suggest that TM loss provides a meaningful signal even in state-space models like Caduceus, although its impact remains more pronounced in transformer-based architectures such as CARMANIA.
Table 8: Effect of TM Loss on Caduceus Fine-Tuned on the Scorpio-Gene–Taxa Dataset. Performance (accuracy, %) across hierarchical taxonomic levels. $\Delta$ denotes the relative improvement when adding TM loss.
Test Set Gene-Heldout Taxa-Heldout
Model Phylum Class Order Family Gene Phylum Class Order Family Gene
$\Delta$ (TM - w/o TM) +6.2 +7.0 +5.8 +4.0 +12.4 +0.3 +0.1 0.1-0.1 0.0 +6.6
Caduceus w/o TM 40.0 24.9 13.4 7.7 26.2 30.3 13.6 4.1 1.5 15.9
Caduceus + TM 46.2 31.9 19.2 11.7 38.6 30.6 13.7 4.0 1.5 22.5
CARMANIA 86.1 76.8 59.6 41.9 90.9 46.9 31.6 15.9 9.4 72.8
A.4 Sensitivity Analysis: Effect of $\beta$
@@
A.5 Higher-Order Transition Matrix
We explored the impact of incorporating TM loss at different Markov orders in CARMANIA, with results shown in Tables 10 and 11. Surprisingly, while the first-order TM consistently improves downstream classification performance across tasks, the second-order TM performs worse than both the first-order TM and the baseline without TM.
@@ A.5 Higher-Order Transition Matrix
A.6 Wide vs. Deep: Architectural Trade-offs for DNA Modeling
@@
A.7 Model Size Scalability
@@
A.8 Impact of Sequence Length Scaling
@@
A.9 Comparison with Other Convolution-Based Models (ConvNova)
To contextualize CARMANIA’s performance with recent convolution-based models, we compared it against ConvNova Bo et al. [2025], a recently proposed architecture for long DNA modeling. Following their evaluation protocol, we used five random seeds on the same test benchmarks to ensure consistency. As shown in Table 16, CARMANIA achieves comparable or superior results on several tasks, demonstrating that its attention-based and Markovian design effectively captures long-range dependencies.
Table 16: Genomics Benchmark Results. Top-1 accuracy (↑) is reported for pretrained HyenaDNA, Caduceus-Ph, ConvNova, and CARMANIA, alongside the CNN baseline. The best result is in bold, and the second-best is underlined.
Task CNN HyenaDNA Caduceus-Ph ConvNova CARMANIA
Enhancers
     Mouse Enhancers 0.730 ± 0.032 0.779 ± 0.013 0.754 ± 0.074 0.784 ± 0.009 0.795 ± 0.002
     Human Enhancers (Cohn) 0.702 ± 0.021 0.718 ± 0.008 0.747 ± 0.004 0.743 ± 0.005 0.725 ± 0.002
     Human Enhancers (Ensembl) 0.744 ± 0.122 0.832 ± 0.006 0.893 ± 0.008 0.900 ± 0.004 0.895 ± 0.002
Species Classification
     Coding vs. Intergenic 0.892 ± 0.008 0.904 ± 0.008 0.915 ± 0.003 0.943 ± 0.001 0.930 ± 0.001
     Human vs. Worm 0.942 ± 0.002 0.961 ± 0.002 0.973 ± 0.001 0.967 ± 0.002 0.970 ± 0.005
Regulatory Elements
     Human Regulatory 0.872 ± 0.005 0.862 ± 0.004 0.872 ± 0.011 0.873 ± 0.002 0.894 ± 0.002
     Human Non-TATA Promoters 0.861 ± 0.009 0.887 ± 0.005 0.946 ± 0.007 0.951 ± 0.003 0.965 ± 0.002
     Human OCR (Ensembl) 0.698 ± 0.013 0.744 ± 0.019 0.828 ± 0.006 0.793 ± 0.004 0.778 ± 0.002
A.10 Pre-training Dataset
@@ A.11 Downstream Datasets
Scorpio-Gene-Taxa Refahi et al. [2025] 10 Multi-class (497–1,929) 114–13,227
Antimicrobial Resistance Prediction Yoo et al. [2024] 3 Multi-class 211–5,274